OpenStudy (anonymous):

Divide. Assume that all expressions are defined. x+3/x^2-2x+1 divided by x+3/x-1 1/x-1 x-1 x^2+6x+9/x-1 1/x^2+1

OpenStudy (anonymous):

\[x+3/x^2-2x+1 .x-1/x+3\]

OpenStudy (anonymous):

if u recirocal x+3/x-1 the sign will change into multiply between the eq.

OpenStudy (anonymous):

x+3 will get cancelled u will be left with x^2-2x+1/x-1

OpenStudy (anonymous):

u understood this much

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{x+3}{x^2-2x+1} = \frac{x+3}{(x-1)(x-1)}\]

OpenStudy (anonymous):

\[\frac{x+3}{(x-1)(x-1)} \times \frac{x-1}{x+3} = \frac{1}{x-1}\]